### v2

# ScholarGym: Benchmarking Large Language Model Capabilities in the Information-Gathering Stage of Deep Research

Abstract
Large language models have advanced from single-turn question answering to deep research systems that iteratively decompose research questions, invoke retrieval tools, and synthesize information across multiple rounds. Evaluating such systems typically involves scoring their final research reports holistically, but this end-to-end paradigm tightly couples the language model’s decision-making, workflow design, and environmental feedback, precluding decomposable analysis of individual components.
[…]
Machine Learning, ICML
1 Introduction
In recent years, large language models (LLMs) have evolved from single-turn question answering to deep research systems capable of long-horizon, iterative information gathering and synthesis (Li et al., 2025c; Wu et al., 2025; Zhang et al., 2025a). In academic literature retrieval, such systems progressively decompose a research question into actionable subqueries, repeatedly invoke retrieval tools, and dynamically adjust their search and filtering strategies through multi-round feedback, ultimately integrating relevant findings into structured research reports or survey-style analyses (Wang et al., 2024; Bao et al., 2025; He et al., 2025). Commercial systems operating in this paradigm—including OpenAI Deep Research, Gemini Deep Research, and Kimi Agent (OpenAI, 2025a; Google, 2025a; Kimi, 2025)—have demonstrated substantial practical value.
As deep research systems advance rapidly, numerous efforts have attempted to evaluate this class of tasks, typically by running a complete deep research pipeline and scoring the final research report holistically (Gou et al., 2025; Du et al., 2025; Li et al., 2025b). While such evaluations provide useful reference points for overall system performance under particular configurations, their principal limitation is that multiple critical factors—the language model’s decision-making, tool interface design, workflow orchestration, and environmental feedback—remain tightly coupled throughout evaluation. These factors mutually influence and constrain one another, making it difficult to attribute observed outcomes to specific components. Under this setup, evaluation results can only indicate whether the system works as a whole, but cannot answer which components are effective and which decisions truly determine research quality.
A principled evaluation framework for deep research should satisfy several key requirements. First, evaluation should center on the language model’s decisions and actions, enabling the contribution of each component to be explicitly disentangled and analyzed, rather than relying solely on end-to-end judgment. Second, since final reports vary widely in structure and style, evaluation should target the upstream behaviors that most determine report quality—information gathering, coverage, and filtering—that have a decisive impact on research quality. Third, to support reproducible and comparable analysis, the evaluation environment should eliminate the non-determinism introduced by live web APIs and open-ended search, characterizing model and component interactions under controlled conditions.
We introduce ScholarGym, an evaluation environment that targets the information-gathering stage of deep research on academic literature. ScholarGymfixes a unified workflow and decomposes it into three stages—Query Planning, Tool Invocation, and Relevance Assessment—with the language model as the sole decision-maker at each stage. All retrieval operates over a static corpus of 570K papers with deterministic ranking, enabling systematic comparison of model behavior during information gathering. We construct 2,536 queries with expert-annotated ground truth and equip the workflow with a memory mechanism that maintains reasoning state across iterations, enabling rigorous, reproducible evaluation of long-horizon academic retrieval.
[…]
2 Related Work
2.1 Deep Research Evaluation
Benchmarks for deep research have evolved from factoid QA to multi-step information synthesis. GAIA (Mialon et al., 2023) and BrowseComp (Wei et al., 2025) evaluate tool use and web navigation but focus on intermediate reasoning rather than comprehensive outputs. DeepResearch Bench (Du et al., 2025; Li et al., 2026), Rigorous Bench (Yao et al., 2025), DEER (Han et al., 2025), and FINDER (Zhang et al., 2025b) shift toward holistic evaluation of report quality using multidimensional rubrics.
A shared limitation is reliance on live APIs. Chen et al. (2025) observe that temporal drift in web content creates a “reproducibility gap” obscuring algorithmic progress. Zhang et al. (2025b) report that evaluation failures frequently stem from anti-scraping mechanisms and URL timeouts rather than model deficiencies. LiveResearchBench (Wang et al., 2025) acknowledges that the absence of fixed ground truth renders verification costly and noisy. These observations motivate controlled evaluation that decouples workflow capabilities from environmental variance.
2.2 Academic Literature Retrieval
Retrieving scientific literature introduces additional challenges: citation networks are dense, terminology is specialized, and relevance judgments require domain expertise. PaSa (He et al., 2025) proposes an autonomous agent for paper search but relies on Google Search APIs, inheriting the non-determinism of general web agents. CiteME (Press et al., 2024) shows that API inconsistencies and broken links cause state-of-the-art models to fail on citation tasks due to environmental friction alone.
Static benchmarks offer a partial remedy. LitSearch (Ajith et al., 2024) constructs a fixed corpus from ACL and ICLR papers and finds that dense retrievers outperform commercial search engines on deep understanding queries. SciNetBench (Shao et al., 2025) uses an OpenAlex snapshot for relation-aware retrieval. HiSciBench (Zhang et al., 2025c) highlights that literature review generation suffers from factuality gaps exacerbated by dynamic sources. Our work extends this direction by coupling a static corpus with an iterative workflow, enabling reproducible assessment of multi-turn planning and selection.
2.3 Deep Research Systems
Current deep research systems fall into two categories. Workflow-based systems decompose queries through explicit planning stages. Alita (Qiu et al., 2025) enables scalable agentic reasoning with minimal predefinition. OWL (Hu et al., 2025b) optimizes multi-agent coordination for task automation. WisPaper (Ju et al., 2025) targets academic search via structured query expansion. These frameworks expose workflow logic but depend on live web APIs, causing execution paths to diverge across runs.
RL-based systems train agents end-to-end on search trajectories. DeepResearcher (Zheng et al., 2025) applies reinforcement learning in real-world environments. Search-R1 (Jin et al., 2025) and SimpleTIR (Xue et al., 2025) train on multi-turn tool interactions. DeepDive (Lu et al., 2025) integrates knowledge graphs with multi-turn RL. Gao et al. (2025) and Li et al. (2025a) scale asynchronous RL for long-horizon search. Tongyi DeepResearch (Team et al., 2025) validates algorithms on a Wikipedia snapshot to address non-stationarity—a “wind tunnel” strategy that parallels our motivation. Step-DeepResearch (Hu et al., 2025a) synthesizes atomic capabilities to bypass unstable API feedback.
Both paradigms face the same obstacle: live web APIs inject distributional drift that confounds learning signals (Xue et al., 2025). ScholarGymprovides a static environment where workflow design and RL training can proceed without environmental noise, enabling controlled comparison across systems.
3 ScholarGym
Figure 1: Overview of ScholarGym. Each iteration comprises three stages connected by directed information flow: solid red arrows indicate stage inputs, and dashed orange arrows indicate outputs. Query Planning receives the subquery tree $\mathcal{M}_{t-1}$, experience buffer $\mathcal{B}_{t-1}$, and feedback $\mathcal{O}_{t-1}$ from the previous iteration, then outputs newly generated child nodes and an updated experience buffer. The resulting subqueries are passed to Tool Invocation, which parameterizes retrieval calls and executes them against the corpus; retrieved candidates undergo relevance-based ranking before selection. Relevance Assessment evaluates ranked papers and produces feedback to guide the next iteration.
We present ScholarGym, a simulation environment for reproducible evaluation of deep research workflows. The environment consists of two components: (1) an iterative deep research workflow with closed-loop feedback, and (2) a benchmark suite containing expert-annotated queries and a unified paper corpus. Figure 1 illustrates the architecture; Table 1 summarizes notation.
Table 1: Notation. Subscript $t$ denotes the iteration index.
Symbol Description
$q$, $d$, $\mathcal{D}$ Research query, date constraint, paper corpus
$\mathcal{G}$ Ground-truth paper set
$\mathcal{M}_{t}$ Subquery tree at iteration $t$
$\mathcal{B}_{t}$ Experience buffer at iteration $t$
$\mathcal{P}_{t}$ Search plan at iteration $t$
$\mathcal{Q}_{t}$ Parameterized tool calls at iteration $t$
$\mathcal{C}_{i}$ Candidate set from $i$-th tool call
$\mathcal{S}_{t}$, $\mathcal{O}_{t}$ Selected papers, feedback at iteration $t$
3.1 Deep Research Workflow
Given a research query $q$, the workflow decomposes it into subqueries—specialized search terms targeting specific aspects of qq—and iteratively expands this set based on retrieval outcomes. Each iteration comprises three stages: (1) query planning analyzes search progress and proposes new subqueries; (2) tool invocation generates and executes parameterized retrieval calls; and (3) relevance assessment evaluates candidate papers and provides feedback $\mathcal{O}_{t}$ to guide the next iteration. The process terminates when no new subqueries are proposed or after $T$ iterations. Algorithm 1 presents the procedure.
Algorithm 1 Deep Research Workflow
1:  Input: Query $q$, date constraint $d$, corpus $\mathcal{D}$, max iterations $T$
2:  Output: Selected paper set $\mathcal{S}$
3:  // Initialization
4:  $\mathcal{M}_{0}\leftarrow\{q\}$, $\mathcal{B}_{0}\leftarrow\emptyset$, $\mathcal{O}_{0}\leftarrow\emptyset$, $\mathcal{S}\leftarrow\emptyset$
5:  for $t=1$ to $T$ do
6:  // Stage 1: Query Planning – analyze progress and propose subqueries
7:  $(\mathcal{P}_{t},\mathcal{M}_{t},\mathcal{B}_{t})\leftarrow\textsc{Plan}(q,\mathcal{M}_{t-1},\mathcal{B}_{t-1},\mathcal{O}_{t-1})$
8:  if $\mathcal{P}_{t}=\emptyset$ then
9:   return $\mathcal{S}$ // no new subqueries, search converged
10:  end if
11:  // Stage 2: Tool Invocation – generate and execute tool calls
12:  $\mathcal{Q}_{t}\leftarrow\textsc{Parameterize}(\mathcal{P}_{t})$
13:  $\{\mathcal{C}_{i}\}_{i=1}^{|\mathcal{Q}_{t}|}\leftarrow\textsc{Execute}(\mathcal{Q}_{t},d,\mathcal{D})$
14:  // Stage 3: Relevance Assessment – evaluate and select papers
15:  $(\mathcal{S}_{t},\mathcal{O}_{t})\leftarrow\textsc{Assess}(\{\mathcal{C}_{i}\},\mathcal{P}_{t})$
16:  $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$
17:  end for
18:  return $\mathcal{S}$
Query Planning.
At iteration $t$, this stage receives the subquery tree $\mathcal{M}_{t-1}$, experience buffer $\mathcal{B}_{t-1}$, and feedback $\mathcal{O}_{t-1}$ from the previous iteration. It identifies underexplored aspects of the research question and produces a plan $\mathcal{P}_{t}$ specifying new subqueries. The subquery tree is rooted at the original query $q$, where each node represents a subquery derived from its parent through semantic refinement.
New subqueries are proposed through three actions: Derive creates a more specific subquery from an existing one (e.g., “transformer efficiency” $\rightarrow$ “sparse attention”); Expand creates a sibling subquery exploring a different aspect; Continue requests additional results for an existing subquery. The plan $\mathcal{P}_{t}$ contains the proposed subqueries with semantic descriptions and relevance criteria to guide subsequent assessment.
Tool Invocation.
This stage receives plan $\mathcal{P}_{t}$ and generates parameterized tool calls $\mathcal{Q}_{t}$. For each subquery in $\mathcal{P}_{t}$, it formulates a structured call $(q_{i},k_{i},d_{i})$ specifying the query string $q_{i}$, number of results $k_{i}$, and date constraint $d_{i}$.
Tool invocation comprises two sub-stages following standard retrieval practice. Retrieval executes calls $\mathcal{Q}_{t}$ against corpus $\mathcal{D}$ using either sparse or dense methods, returning an initial candidate set for each query. Sparse methods excel at exact term matching, while dense methods capture semantic relationships beyond lexical overlap. Ranking reorders candidates within each set by relevance scores derived from the underlying retrieval method. The ranked sets $\{\mathcal{C}_{i}\}$ are passed to relevance assessment, where each $\mathcal{C}_{i}$ contains up to $k_{i}$ papers with metadata (title, abstract, publication date, identifier). Implementation details appear in Appendix C.2.
Relevance Assessment.
This stage evaluates each candidate paper against the research objective using the criteria in $\mathcal{P}_{t}$. We consider two strategies: Abstract-only classification determines relevance from titles and abstracts; Adaptive browsing permits an “uncertain” label for ambiguous cases, triggering full-text examination before final classification. The stage outputs selected papers $\mathcal{S}_{t}$ and feedback $\mathcal{O}_{t}$, which summarizes the current search outcomes and suggests refinements for the next iteration.
Memory Mechanism.
Long-horizon search requires coherent state across iterations. We design two structures: the subquery tree $\mathcal{M}_{t}$ organizes subqueries hierarchically, recording derivation paths and retrieved papers for each node; the experience buffer $\mathcal{B}_{t}$ compresses search history into a fixed-length summary, preventing context overflow while preserving key insights.
3.2 Benchmark Construction
We construct ScholarGymfrom two established academic retrieval datasets: PaSa (He et al., 2025) and LitSearch (Ajith et al., 2024). Queries originate from three splits: PaSa-AutoScholar (generated from citation contexts), PaSa-RealScholar (human-curated research questions), and LitSearch (real-world literature search scenarios). Each query is paired with expert-annotated ground-truth papers. For LitSearch queries lacking explicit temporal bounds, we assign date constraints matching the latest publication date among ground-truth papers. Papers from both datasets are aggregated, deduplicated by arXiv identifier, and enriched with metadata via the arXiv API, yielding a corpus $\mathcal{D}$ of 570K papers spanning computer science, physics, and mathematics.
Evaluation Subsets.
The benchmark is partitioned into two subsets. Test-Fast contains 200 queries sampled for balanced coverage across sources, enabling rapid iteration during development. Test-Hard contains 100 queries on which all evaluated models perform poorly; these queries tend to have larger ground-truth sets (average 2.6 papers) and require finding papers across multiple research areas. Table 2 summarizes dataset statistics; Appendix D provides detailed construction procedures.
Table 2: Dataset statistics. #GT denotes average ground-truth papers per query; Len. denotes query length in characters.
Subset #Query Avg. #GT Avg. Len.
Test-Fast 200 1.9 113.0
Test-Hard 100 2.6 101.8
ALL 2,536 2.3 110.4
4 Experiments
4.1 Experimental Setup
Backbone Models.
We evaluate both open-source and proprietary LLMs. Open-source models include Qwen3 (Yang et al., 2025) (8B, 30B) and GLM-4.7 (Zeng et al., 2025). Proprietary models include DeepSeek-V3.2 (Liu et al., 2025), GPT-5.2 (OpenAI, 2025b), and Gemini3-Pro (Google, 2025b). For models that support optional extended thinking (Qwen3 and DeepSeek-V3.2), we evaluate both standard and thinking-enabled configurations; the latter are denoted with $\dagger$.
Baselines.
We compare against a Direct Query baseline that bypasses iterative query planning entirely. Under this baseline, the original research query $q$ is submitted once to the retrieval backend without decomposition; retrieved candidates then undergo the same relevance assessment procedure as the full workflow.
Retrieval Backend.
We implement two retrieval methods over paper titles and abstracts. (1) Sparse retrieval employs BM25, which scores documents by term frequency-inverse document frequency weighting and excels at exact lexical matching. (2) Dense retrieval encodes documents using Qwen3-Embedding-0.6B and performs approximate nearest neighbor search in a Qdrant vector database, capturing semantic similarity beyond surface-level term overlap. Unless otherwise noted, sparse retrieval serves as the default backend.
Evaluation Metrics.
Let $\mathcal{G}$ denote the ground-truth set, $\mathcal{R}$ the retrieved candidates, and $\mathcal{S}$ the final selected papers. We report metrics at two stages: retrieval (prefixed “Ret.”) measures candidate quality before filtering; selection evaluates end-to-end performance. Recall and precision are defined as:
$\text{R}=\frac{|\mathcal{S}\cap\mathcal{G}|}{|\mathcal{G}|},\quad\text{P}=\frac{|\mathcal{S}\cap\mathcal{G}|}{|\mathcal{S}|}$ (1)
$\text{Ret.R}=\frac{|\mathcal{R}\cap\mathcal{G}|}{|\mathcal{G}|},\quad\text{Ret.P}=\frac{|\mathcal{R}\cap\mathcal{G}|}{|\mathcal{R}|}$ (2)
F1 is the harmonic mean of R and P. We additionally introduce two diagnostic metrics:
  • Avg.Distance quantifies query planning quality. For each ground-truth paper $g\in\mathcal{G}$, let $r_{g}$ denote its best rank across all subqueries. With cutoff c=100c{=}100:
    $\text{Avg.Dist}=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\max\left(1-\frac{r_{g}}{c},0\right)$ (3)
    Higher values indicate relevant papers are surfaced earlier in retrieval results.
  • GT Discard Rate measures assessment errors—the fraction of discarded candidates that were actually relevant:
    $\text{GT Disc.}=\frac{|(\mathcal{R}\cap\mathcal{G})\setminus\mathcal{S}|}{|\mathcal{R}\setminus\mathcal{S}|}$ (4)
Implementation.
We use Test-Fast with T=5T{=}5 iterations and greedy decoding (Temperature=0, Top_p=1) for reproducibility. Main results (Table 3) report metrics at the final iteration under sparse retrieval and Abstract-only assessment mode.
4.2 Main Results
Table 3: Performance on Test-Fast and Test-Hard benchmarks (sparse retrieval, Abstract-only, iteration 5). R/P/F1 denote selection-stage recall, precision, and F1; Ret.R/Ret.P/Ret.F1 denote retrieval-stage metrics. Best results are bolded; second-best are underlined. $\dagger$ indicates extended thinking mode.
Test-Fast Test-Hard
Model R P F1 Ret.R Ret.P Ret.F1 R P F1 Ret.R Ret.P Ret.F1
Direct Query Baseline
Qwen3-8B 0.185 0.042 0.069 0.210 0.036 0.061 0.048 0.004 0.007 0.055 0.003 0.006
Qwen3-30B 0.312 0.058 0.098 0.350 0.052 0.091 0.089 0.006 0.011 0.095 0.005 0.010
Open-Source Models
Qwen3-8B 0.483 0.152 0.231 0.550 0.022 0.042 0.123 0.016 0.028 0.165 0.005 0.010
Qwen3-8B 0.458 0.216 0.293 0.580 0.023 0.045 0.120 0.013 0.023 0.213 0.007 0.014
Qwen3-30B 0.673 0.181 0.285 0.720 0.011 0.021 0.226 0.016 0.030 0.316 0.006 0.011
Qwen3-30B 0.482 0.290 0.362 0.590 0.025 0.048 0.172 0.058 0.087 0.264 0.010 0.019
GLM-4.7 0.754 0.111 0.194 0.814 0.010 0.020 0.216 0.017 0.031 0.398 0.006 0.012
Proprietary Models
DeepSeek-V3.2 0.855 0.135 0.233 0.862 0.007 0.014 0.372 0.017 0.032 0.430 0.006 0.012
DeepSeek-V3.2 0.812 0.287 0.423 0.872 0.009 0.018 0.355 0.041 0.074 0.448 0.008 0.016
GPT-5.2 0.837 0.305 0.447 0.883 0.009 0.018 0.397 0.045 0.081 0.505 0.008 0.016
Gemini3-Pro 0.950 0.199 0.329 0.958 0.011 0.022 0.365 0.028 0.052 0.504 0.010 0.020
Table 3 presents performance on both Test-Fast and Test-Hard benchmarks.
Iterative Planning Effectiveness.
[…]
Overall Performance.
GPT-5.2 achieves the highest F1 (0.447) by balancing recall (0.837) and precision (0.305). Gemini3-Pro attains the highest recall (0.950) but at lower precision (0.199), yielding F1 of 0.329. A substantial performance gap separates proprietary and open-source models: the best open-source result (Qwen3-30B, F1=0.362) trails GPT-5.2 by 19% in relative terms. We analyze the sources of this gap in Section 4.4.
Retrieval vs. Selection Performance.
Comparing Ret.R and R reveals where ground-truth coverage is lost during assessment. Gemini3-Pro retrieves 95.8% of ground-truth papers and retains 95.0% after assessment—a gap of only 0.8 percentage points. In contrast, Qwen3-8B exhibits a 12.2 point drop (Ret.R=0.580 to R=0.458) due to aggressive filtering. This gap is inversely correlated with model capability, indicating that relevance assessment represents a high-leverage optimization target for smaller models.
Extended Thinking Trade-offs.
Extended thinking induces a precision-recall trade-off. Qwen3-8B achieves 42% higher precision than its base variant (0.216 vs. 0.152) while sacrificing 5% recall (0.458 vs. 0.483). This trade-off arises from more aggressive filtering during relevance assessment, which we analyze in Section 4.4.2. The net F1 improvement scales with model capability: DeepSeek-V3.2 gains +82% F1 over DeepSeek-V3.2 (0.423 vs. 0.233), whereas Qwen3-8B gains +27% (0.293 vs. 0.231).
Model Scaling.
Within the Qwen3 family, scaling from 8B to 30B yields substantial recall improvement (+39%, from 0.483 to 0.673) but only marginal precision gain (0.181 vs. 0.152). Larger models generate more diverse subqueries that cover broader semantic regions, yet do not necessarily impose stricter relevance criteria during assessment.
Generalization to Test-Hard.
All models degrade substantially on Test-Hard, which comprises cross-domain queries with larger ground-truth sets. The best Test-Fast F1 (0.447) drops to 0.087—an 80% decline. Notably, Qwen3-30B achieves the highest Test-Hard F1 (0.087), marginally outperforming GPT-5.2 (0.081). However, this advantage derives from higher precision (0.058 vs. 0.045) rather than recall; GPT-5.2 maintains the highest Test-Hard recall (0.397 vs. 0.172). The relative F1 gain of thinking-enabled models on cross-disciplinary queries reflects their conservative selection strategy rather than improved discovery coverage.
[…]
4.3 Iteration Dynamics
Table 4: Per-iteration recall and precision on Test-Fast (sparse retrieval, Abstract-only).
Model It.1 It.2 It.3 It.4 It.5
Recall
Qwen3-8B 0.332 0.403 0.440 0.453 0.483
Qwen3-8B 0.298 0.400 0.435 0.450 0.458
Qwen3-30B 0.526 0.626 0.659 0.669 0.673
Qwen3-30B 0.375 0.410 0.451 0.460 0.482
GLM-4.7 0.515 0.602 0.659 0.714 0.754
DeepSeek-V3.2 0.540 0.655 0.745 0.810 0.855
DeepSeek-V3.2 0.498 0.615 0.710 0.775 0.812
GPT-5.2 0.618 0.725 0.772 0.805 0.837
Gemini3-Pro 0.694 0.812 0.882 0.925 0.950
Precision
Qwen3-8B 0.177 0.149 0.144 0.148 0.152
Qwen3-8B 0.218 0.245 0.241 0.227 0.216
Qwen3-30B 0.227 0.195 0.185 0.185 0.181
Qwen3-30B 0.292 0.291 0.295 0.290 0.290
GLM-4.7 0.182 0.158 0.136 0.115 0.111
DeepSeek-V3.2 0.220 0.172 0.148 0.141 0.135
DeepSeek-V3.2 0.395 0.341 0.311 0.297 0.287
GPT-5.2 0.345 0.328 0.315 0.310 0.305
Gemini3-Pro 0.229 0.199 0.183 0.199 0.199
[…]
Table 4 presents per-iteration metrics, while Figure 2 extends the analysis to 25 iterations for detailed convergence characterization.
Convergence Patterns.
Two distinct regimes emerge across model capabilities. High-capability models front-load discovery: Gemini3-Pro achieves 73% of its final recall (0.694/0.950) in iteration 1, with diminishing marginal gains of +12%, +7%, +4%, and +3% in subsequent iterations—reflecting effective initial query decomposition. Lower-capability models exhibit more gradual convergence: Qwen3-8B gains +21%, +9%, +3%, and +7% across iterations 2–5, deriving greater benefit from feedback-driven query refinement.
Precision Dynamics.
Precision decays monotonically as iterations accumulate candidates. DeepSeek-V3.2 decreases from 0.395 to 0.287 (-27%); GLM-4.7 from 0.182 to 0.111 (-39%). A notable exception is Qwen3-30B, which maintains stable precision (0.292 to 0.290) throughout the iteration sequence. This stability suggests that thinking-enabled assessment enforces consistent relevance thresholds irrespective of candidate volume, whereas standard models progressively lower their acceptance criteria.
Saturation Beyond T=5T{=}5.
Figure 2 demonstrates that all Qwen3 variants plateau within 10 iterations. Beyond T=5T{=}5, Qwen3-30B gains only 2.3 percentage points (0.673 to 0.702); Qwen3-8B gains merely 0.4 points (0.483 to 0.487). Thinking-enabled variants exhibit analogous saturation patterns. This ceiling reflects exhaustion of productive query reformulations—later iterations revisit previously explored semantic regions rather than discovering novel relevant work. These observations motivate our default setting of T=5T{=}5 as an effective trade-off between coverage and computational cost.
Key Findings (1) High-capability models achieve 70–90% of final recall within two iterations. (2) Precision decays with iterations for most models; thinking-enabled variants maintain more stable precision. (3) Recall saturates by T=10T{=}10; gains beyond T=5T{=}5 are minimal.
4.4 Component Analysis
4.4.1 Query Planning Quality
Table 5: Avg.Distance by iteration on Test-Fast (sparse retrieval). Higher values indicate ground-truth papers rank earlier in retrieval results.
Model It.1 It.2 It.3 It.4 It.5 Avg.
Qwen3-8B 0.588 0.600 0.604 0.587 0.583 0.592
Qwen3-8B 0.582 0.626 0.626 0.623 0.616 0.615
Qwen3-30B 0.695 0.735 0.731 0.702 0.708 0.714
Qwen3-30B 0.668 0.685 0.678 0.672 0.667 0.674
GLM-4.7 0.730 0.771 0.775 0.798 0.813 0.777
DeepSeek-V3.2 0.739 0.773 0.809 0.835 0.845 0.800
DeepSeek-V3.2 0.737 0.800 0.831 0.820 0.816 0.801
GPT-5.2 0.771 0.803 0.820 0.826 0.836 0.811
Gemini3-Pro 0.784 0.828 0.855 0.876 0.881 0.845
Table 5 presents Avg.Distance trajectories across iterations.
Model Hierarchy.
Gemini3-Pro attains the highest Avg.Distance (0.845), with subqueries ranking ground-truth papers within the top 15–20 positions on average. A substantial gap separates proprietary models (0.800–0.845) from open-source alternatives (0.592–0.777). This 43% disparity between Gemini3-Pro and Qwen3-8B accounts for much of the recall difference observed in Table 3: well-formulated queries surface relevant papers earlier in the ranking, increasing discovery probability within the fixed retrieval budget.
Iteration Dynamics.
Planning quality improves across iterations through feedback incorporation. Gemini3-Pro increases from 0.784 (iteration 1) to 0.881 (iteration 5), representing a 12% gain. Proprietary models exhibit steeper improvement trajectories (+10% on average) compared to open-source counterparts (+3%), suggesting more effective utilization of feedback signals from prior retrieval outcomes.
Key Findings (1) Proprietary models achieve 43% higher Avg.Distance than open-source alternatives. (2) Planning quality improves across iterations via feedback incorporation, with proprietary models showing steeper gains (+10% vs. +3%).
4.4.2 Relevance Assessment
Figure 3: Per-iteration GT Discard Rate (%) on Test-Fast. Darker cells indicate higher discard rates of ground-truth papers during relevance assessment.
Table 6: GT Discard Rate (%) and precision by assessment strategy on Test-Fast.
Abstract-only Adaptive
Model Disc.% Prec. Disc.% Prec.
Qwen3-8B 0.66 0.152 0.34 0.149
Qwen3-8B 0.93 0.216 0.58 0.218
Qwen3-30B 0.13 0.181 0.29 0.217
Qwen3-30B 1.03 0.290 0.71 0.311
GLM-4.7 0.36 0.111 0.23 0.128
DeepSeek-V3.2 0.12 0.135 0.10 0.167
DeepSeek-V3.2 0.17 0.287 0.12 0.302
GPT-5.2 0.14 0.305 0.10 0.320
Gemini3-Pro 0.14 0.199 0.11 0.232
Figure 3 and Table 6 characterize relevance assessment behavior.
Error Patterns.
Proprietary models maintain GT Discard Rates below 0.20% across all iterations (Figure 3), with errors distributed randomly. Thinking-enabled open-source models exhibit elevated, iteration-dependent rates: Qwen3-8B starts at 1.46% in iteration 1 and declines to 0.38% by iteration 5; Qwen3-30B remains consistently elevated (1.22% to 0.77%). This trajectory indicates that extended thinking applies aggressive relevance thresholds that gradually calibrate through accumulated assessment experience.
Precision-Error Trade-off.
Higher GT Discard Rates correlate with elevated precision, quantifying the cost of stringent filtering. Qwen3-30B achieves the highest open-source precision (0.290) alongside the highest discard rate (1.03%), whereas Qwen3-30B maintains minimal discard rates (0.13%) with lower precision (0.181). This pattern explains the precision-recall trade-off observed in Section 4.2: thinking-enabled models improve precision through aggressive filtering at the cost of discarding some relevant papers.
Adaptive Browsing.
Full-text retrieval for ambiguous candidates reduces GT Discard Rates while preserving precision. Qwen3-8B improves from 0.93% to 0.58%; Qwen3-30B from 1.03% to 0.71%. Proprietary models show negligible change, as their abstract-based assessment already achieves high accuracy. The benefit is most pronounced for methodology papers where relevance depends on technical details absent from abstracts.
Key Findings (1) Thinking-enabled models achieve higher precision via aggressive filtering (GT Discard Rates: 0.93–1.03% vs. <<0.20% for proprietary models). (2) Adaptive Browsing reduces discard rates by 30–40% for open-source models.
4.5 Ablation Studies
4.5.1 Retrieval Backend
Table 7: Sparse versus dense retrieval on Test-Fast (Abstract-only, iteration 5). Dense retrieval uses Qwen3-Embedding-0.6B for document encoding.
Sparse Dense
Model R P F1 R P F1
Qwen3-8B 0.483 0.152 0.231 0.608 0.183 0.281
Qwen3-8B 0.458 0.216 0.293 0.488 0.209 0.293
Qwen3-30B 0.673 0.181 0.285 0.718 0.208 0.322
Qwen3-30B 0.482 0.290 0.362 0.532 0.287 0.373
Table 7 compares both sparse and dense retrieval backends.
Sparse vs. Dense Retrieval.
Dense retrieval yields model-dependent recall improvements. Standard models exhibit substantial gains: Qwen3-8B improves by 12.5 percentage points (0.483 to 0.608), and Qwen3-30B by 4.5 points (0.673 to 0.718). Thinking-enabled models show smaller improvements (3.0–5.0 points), likely because their stricter assessment thresholds limit the net recall gain even when additional relevant candidates are surfaced.
Key Findings Dense retrieval improves recall by 7–26% for standard models, with smaller gains (7–10%) for thinking-enabled variants due to stricter assessment filtering.
4.5.2 Memory Mechanism
Table 8: Effect of the memory mechanism on Test-Fast performance. “w/o Memory” removes the experience buffer $\mathcal{B}_{t}$, retaining only raw conversation history as context input to the Query Planner.
With Memory w/o Memory
Model R P F1 R P F1
Qwen3-8B 0.483 0.152 0.231 0.485 0.140 0.217
Qwen3-8B 0.458 0.216 0.293 0.411 0.168 0.239
Qwen3-30B 0.673 0.181 0.285 0.651 0.169 0.268
Qwen3-30B 0.482 0.290 0.362 0.402 0.218 0.282
Table 8 quantifies the contribution of the memory mechanism.
Performance Impact.
The experience buffer $\mathcal{B}_{t}$ aggregates the reasoning traces and outputs from all prior query planning iterations into a compressed summary, tracking discovered papers and explored query regions to guide subsequent iterations toward unexplored areas. Removing this component degrades performance uniformly across models. Thinking-enabled models exhibit the largest drops: Qwen3-8B decreases from 0.293 to 0.239 F1 (-18.4%), and Qwen3-30B from 0.362 to 0.282 (-22.1%). Standard models show more modest degradation (6.0–6.1%).
Compression vs. Raw History.
Without the experience buffer, query planning receives raw conversation history instead of a compressed summary. The lengthier, unstructured context impedes effective extraction of exploration state, leading to regeneration of semantically similar subqueries and wasted retrieval budget. Thinking-enabled models are disproportionately affected: their verbose reasoning traces produce longer histories that exacerbate context utilization difficulties, whereas the compressed buffer distills prior exploration into an efficiently consumable format.
Key Findings The memory mechanism prevents query redundancy across iterations. Removing $\mathcal{B}_{t}$ degrades F1 by 6–22%, with thinking-enabled models most affected.
5 Conclusion
We introduced ScholarGymto address the reproducibility challenge in deep research evaluation. By grounding agentic workflows in a static, deterministic environment containing 570K papers, we isolated algorithmic reasoning from live API variance, enabling precise dissection of model behaviors. Our modular analysis reveals distinct operating regimes: while proprietary models maximize discovery through superior query formulation, “extended thinking” paradigms in open models function primarily as high-precision filters, often at the cost of recall. We further identified query planning quality and relevance assessment calibration as the dual bottlenecks limiting open-source performance on complex tasks.
Our findings position ScholarGymas a critical testbed for next-generation research agents. The deterministic environment specifically enables the training of reinforcement learning policies for query planning without the variance of live web search, paving the way for agents that can learn optimal exploration-exploitation trade-offs in information seeking. We open-source our environment and benchmarks to facilitate this community effort.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experimental Results
A.1 Additional Visualizations
Figure 4 presents detailed recall and precision trajectories across iterations. Figure 5 shows query planning quality trajectories across iterations. Figures 6 and 7 present ablation study results.
Figure 4: Recall and Precision trajectories across 5 iterations on Test-Fast (sparse retrieval, Abstract-only).
Figure 5: Avg.Distance trajectories across iterations. Higher values indicate queries that rank ground-truth papers earlier in retrieval results.
Figure 6: Sparse versus dense retrieval recall comparison.
Figure 7: Impact of the memory mechanism. Removing the experience buffer $\mathcal{B}_{t}$ degrades F1 across all models, with percentage drops annotated.
A.2 Detailed Test-Hard Results
Table 9 presents comprehensive performance metrics on the Test-Hard benchmark. Test-Hard queries span multiple research domains and exhibit higher complexity than Test-Fast, resulting in substantially lower performance across all models.
Table 9: Detailed performance on Test-Hard benchmark (sparse retrieval, Abstract-only, iteration 5). R/P/F1 denote selection-stage metrics; Ret.R/Ret.P/Ret.F1 denote retrieval-stage metrics. Avg.Dist measures query planning effectiveness.
Model Avg.Dist R P F1 Ret.R Ret.P Ret.F1
Open-Source Models
Qwen3-8B 0.270 0.123 0.016 0.028 0.165 0.005 0.010
Qwen3-8B 0.258 0.120 0.013 0.023 0.213 0.007 0.014
Qwen3-30B 0.311 0.226 0.016 0.030 0.316 0.006 0.011
Qwen3-30B 0.264 0.172 0.058 0.087 0.264 0.010 0.019
GLM-4.7 0.341 0.216 0.017 0.031 0.398 0.006 0.012
Proprietary Models
DeepSeek-V3.2 0.387 0.372 0.017 0.032 0.430 0.007 0.014
DeepSeek-V3.2 0.398 0.355 0.041 0.074 0.448 0.008 0.015
GPT-5.2 0.383 0.397 0.045 0.081 0.505 0.008 0.016
Gemini3-Pro 0.404 0.365 0.028 0.052 0.504 0.011 0.020
A.3 Cumulative Metrics on Test-Hard
Table 10 shows per-iteration recall and precision on Test-Hard. Compared to Test-Fast, all models exhibit slower convergence and substantially lower final performance.
Table 10: Cumulative recall and precision at each iteration on Test-Hard (sparse retrieval, Abstract-only).
Recall Precision
Model It.1 It.2 It.3 It.4 It.5 It.1 It.2 It.3 It.4 It.5
Qwen3-8B 0.080 0.091 0.111 0.111 0.123 0.020 0.013 0.019 0.017 0.016
Qwen3-8B 0.035 0.073 0.077 0.104 0.120 0.015 0.016 0.011 0.013 0.013
Qwen3-30B 0.126 0.155 0.204 0.206 0.226 0.023 0.020 0.019 0.016 0.016
Qwen3-30B 0.118 0.143 0.149 0.154 0.172 0.054 0.058 0.047 0.044 0.058
GLM-4.7 0.086 0.113 0.169 0.188 0.216 0.023 0.017 0.017 0.017 0.017
DeepSeek-V3.2 0.235 0.290 0.330 0.355 0.372 0.027 0.022 0.019 0.018 0.017
DeepSeek-V3.2 0.220 0.275 0.315 0.340 0.355 0.056 0.049 0.044 0.042 0.041
GPT-5.2 0.285 0.338 0.365 0.382 0.397 0.052 0.048 0.047 0.046 0.045
Gemini3-Pro 0.265 0.310 0.337 0.352 0.365 0.033 0.028 0.026 0.028 0.028
A.4 Impact of Adaptive Browsing on Test-Hard
Table 11 compares Abstract-only and Adaptive Browsing strategies on Test-Hard. While Adaptive Browsing improves recall for most models, the precision gains are less consistent than on Test-Fast.
Table 11: Abstract-only versus Adaptive Browsing on Test-Hard (sparse retrieval, iteration 5).
Abstract-only Adaptive
Model R P F1 R P F1
Qwen3-8B 0.123 0.016 0.028 0.188 0.026 0.046
Qwen3-8B 0.120 0.013 0.023 0.179 0.019 0.035
Qwen3-30B 0.226 0.016 0.030 0.256 0.035 0.061
Qwen3-30B 0.172 0.058 0.087 0.160 0.042 0.066
GLM-4.7 0.216 0.017 0.031 0.317 0.023 0.042
DeepSeek-V3.2 0.372 0.017 0.032 0.385 0.021 0.040
DeepSeek-V3.2 0.355 0.041 0.074 0.368 0.043 0.078
GPT-5.2 0.397 0.045 0.081 0.422 0.050 0.090
Gemini3-Pro 0.365 0.028 0.052 0.372 0.033 0.062
Appendix B Evaluation Metrics
We evaluate workflow performance at two stages: retrieval (candidates returned by the search backend) and selection (papers retained after relevance assessment). Let $\mathcal{G}$ denote the ground-truth paper set, $\mathcal{R}_{t}$ the papers retrieved at iteration $t$, and $\mathcal{S}_{t}$ the papers selected at iteration $t$. Cumulative sets are defined as $\mathcal{R}=\bigcup_{t=1}^{T}\mathcal{R}_{t}$ and $\mathcal{S}=\bigcup_{t=1}^{T}\mathcal{S}_{t}$.
Retrieval-Stage Metrics.
Retrieval precision and recall measure the quality of candidates before relevance filtering:
Ret.Precision $=\frac{|\mathcal{R}\cap\mathcal{G}|}{|\mathcal{R}|},\quad\text{Ret.Recall}=\frac{|\mathcal{R}\cap\mathcal{G}|}{|\mathcal{G}|}.$ (5)
Ret.F1 is the harmonic mean of Ret.Precision and Ret.Recall.
Selection-Stage Metrics.
Selection precision and recall evaluate the final output:
Precision $=\frac{|\mathcal{S}\cap\mathcal{G}|}{|\mathcal{S}|},\quad\text{Recall}=\frac{|\mathcal{S}\cap\mathcal{G}|}{|\mathcal{G}|}.$ (6)
F1 is the harmonic mean of Precision and Recall.
Average Distance.
This metric quantifies query planning effectiveness by measuring how early ground-truth papers appear in retrieval rankings. For each ground-truth paper $g\in\mathcal{G}$, let $r_{g}$ denote its minimum rank across all subqueries and iterations. Given a cutoff $c$ (default 100), the distance score is:
$\text{dist}(g)=\max\left(1-\frac{r_{g}}{c},0\right).$ (7)
Papers ranked beyond position $c$ receive zero credit. Avg.Distance averages this score:
$\text{Avg.Distance}=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\text{dist}(g).$ (8)
Higher values indicate that generated subqueries successfully surface ground-truth papers near the top of retrieval results.
Ground-Truth Discard Rate.
This diagnostic metric measures relevance assessment errors—ground-truth papers that were retrieved but subsequently discarded:
$\text{GT Discard Rate}=\frac{|(\mathcal{R}\cap\mathcal{G})-(\mathcal{S}\cap\mathcal{G})|}{|\mathcal{R}\cap\mathcal{G}|}.$ (9)
Lower values indicate better retention of relevant papers during assessment.
Per-Iteration Metrics.
For iteration-level analysis, we compute cumulative metrics at each step. Let $\mathcal{R}^{(t)}=\bigcup_{i=1}^{t}\mathcal{R}_{i}$ and $\mathcal{S}^{(t)}=\bigcup_{i=1}^{t}\mathcal{S}_{i}$ denote cumulative sets through iteration $t$. Per-iteration metrics substitute these sets in the formulas above, enabling analysis of convergence patterns and marginal gains.
Appendix C Implementation Details
C.1 Query Planning
The query planning stage generates subqueries through structured reasoning over the current memory state. This stage maintains a subquery tree rooted at the original query (id=0) and validates consistency between requested and actual retrieval counts across iterations.
Subquery Generation Principles.
Seven principles guide the planning process: (1) preserve key concepts from the original query; (2) avoid domain-specific terms unless explicitly relevant; (3) employ synonyms for comprehensive coverage; (4) vary phrasing to maximize diversity; (5) prefer general expressions over narrow descriptions; (6) include standard technical terminology; (7) balance precision and recall.
Subquery Operations.
Three operations govern subquery generation:
  • Continue: Request additional results for an existing subquery via pagination.
  • Derive: Create a specialized subquery that narrows the search scope from an existing node.
  • Expand: Create a parallel subquery exploring orthogonal aspects at the same granularity.
Memory Components.
Two memory structures are maintained: the experience buffer provides cumulative long-term memory across iterations; the checklist specifies concrete retrieval criteria for each subquery to guide relevance assessment.
Output Schema.
The query planning stage outputs structured JSON containing subqueries (with link type, source id, text, and target count), a checklist, an experience replay summary, and a completion flag.
C.2 Tool Invocation
The tool invocation stage executes retrieval calls against the static paper corpus $\mathcal{D}$. We implement two retrieval methods:
Dense Vector Retrieval.
Papers are embedded using Qwen3-Embedding-0.6B on concatenated title and abstract text. Embeddings are stored in a Qdrant vector database. Given a subquery, we compute its embedding and retrieve candidates by cosine similarity. Date filtering ensures candidates satisfy temporal constraints.
Sparse Retrieval.
We build a BM25 index over tokenized title and abstract text. Queries are tokenized identically, and candidates are ranked by BM25 scores. This method excels at matching specific technical terms that may be underweighted in dense representations.
Both methods support pagination via Continue operations. The static corpus ensures identical calls yield identical results across runs.
C.3 Relevance Assessment
Relevance assessment evaluates candidates against the research objective. This stage prioritizes: direct topical alignment, clear methodological contributions, recent or seminal works, strong empirical evidence, and complete abstracts. Survey papers are down-ranked unless explicitly requested.
Assessment Modes.
Two modes are implemented:
Abstract-only mode classifies papers as selected or discarded based on titles and abstracts. This mode prioritizes efficiency.
Adaptive browsing mode permits an “uncertain” classification for ambiguous cases. Papers marked to_browse undergo full-text extraction, after which the relevance assessment stage re-evaluates with augmented information.
Feedback.
The assessment produces structured feedback: topics covered, synopsis of selected and discarded papers, and suggestions for query refinement. This propagates to query planning in subsequent iterations.
C.4 Browser Module
The browser module examines full-text content under adaptive browsing mode. Given a paper and an extraction goal, it fetches HTML content from ar5iv, parses hierarchical sections with mathematical notation preservation, and extracts targeted information.
The ar5iv parser extracts title, authors, abstract, and hierarchical sections. Stop-word sections (references, acknowledgments) are filtered. Mathematical expressions are preserved from ltx_Math elements. The extraction produces: relevant section names, verbatim excerpts, and a synthesis of at most three sentences.
Appendix D Benchmark Construction
Source Processing.
From PaSa (He et al., 2025), we extract AutoScholar (citation-derived queries) and RealScholar (human-curated queries) test sets. From LitSearch (Ajith et al., 2024), we map Semantic Scholar identifiers to arXiv IDs using external ID fields and regex extraction from PDF URLs.
Corpus Construction.
Papers are merged by arXiv identifier. We query the arXiv API for complete abstracts, standardized dates, author lists, and category labels. Papers with empty abstracts, invalid identifiers, or dates outside 1990–2024 are excluded. The final corpus contains 570K papers spanning computer science, physics, and mathematics.
Index Construction.
For dense retrieval, we embed papers with Qwen3-Embedding-0.6B and store vectors in Qdrant. For sparse retrieval, we build a BM25 index over tokenized text. Both indices are static, ensuring reproducibility.
Subset Selection.
Test-Fast samples 200 queries with proportional source representation and ground-truth diversity. Test-Hard selects 100 challenging queries with larger ground-truth sets and cross-domain requirements.
Appendix E Prompt Templates
This section presents the prompt templates for each stage of our deep research workflow. All prompts adopt a structured XML-style input format and require JSON-formatted outputs for robust parsing.
E.1 Query Planning Prompt
This stage receives the subquery tree $\mathcal{M}_{t-1}$, experience buffer $\mathcal{B}_{t-1}$, and feedback $\mathcal{O}_{t-1}$ to produce plan $\mathcal{P}_{t}$ (cf. Section 3.1). The prompt encodes three operations—Continue, Derive, and Expand—along with subquery generation strategies. Table 12 details the prompt components; Figure 8 presents the template; Figure 9 shows a representative output.
Table 12: Query Planning prompt components: strategies, linking guidance, and field reference.
Panel A: Strategies (Subquery Generation Principles)
S1 Preserve all key points from the user’s original query.
S2 Avoid domain-specific keywords unless explicitly requested or strongly domain-tied.
S3 Use synonyms and near-synonyms to capture different expressions of the same concept.
S4 Avoid repetitive sentence structures; vary phrasing to improve diversity.
S5 Prefer general, widely-used expressions over narrow descriptions.
S6 Include professional and technical terms commonly used in the field.
S7 Combine multiple strategies to generate queries that are both precise and diverse.
Panel B: Linking Guidance (Subquery Operations)
Continue Request additional results for an existing subquery via pagination.
Required: source_id, target_k.
Derive Create a specialized subquery that deepens exploration from an existing node.
Required: source_id, text, target_k.
Expand Create a parallel subquery exploring alternative aspects at the same granularity.
Required: source_id, text, target_k.
Rules The subquery space is a tree rooted at id=0. Every source_id must reference a node in <all_subqueries>. Root (id=0) is valid for Derive/Expand but not for Continue.
Panel C: Field Reference (Input Fields)
all_subqueries Each line: id, source (parent id), link type, iteration, text.
last_iteration_state Per subquery: id, text, target_k, retrieved, selected, overview.
previous_iteration_state Earlier iteration states, same structure as above.
last_checklist Concrete retrieval criteria from the previous round.
last_experience_replay Cumulative long-term memory across iterations.
Prompt: Query Planning [System]
You are the Query Planning module in a Deep Research workflow. Analyze the user query and the research memory to propose subqueries and set target_k (papers per subquery). The subquery space is a tree rooted at id=0 (the original user query). You may choose any existing node (including id=0) as the source for DERIVE/EXPAND, and any non-root existing node for CONTINUE. Validate the last_iteration_state against the last_checklist to ensure consistency (e.g., whether requested target_k and actually retrieved/selected counts align). The Relevance Assessment module provides overview for all subqueries from all previous rounds. The most recent overviews can be found in last_iteration_state, while earlier ones are stored in previous_iteration_state. Use experience_replay as the cumulative long-term memory: integrate last_experience_replay, insights from last_iteration_state, and your current planning rationale; emphasize what changed since the previous iteration. Maintain a concise checklist of concrete retrieval criteria per subquery (what to retrieve next: required methods, tasks, datasets, time ranges, include/exclude signals) for the Relevance Assessment module and the next Query Planning iteration. Each retrieval call returns exactly one page with up to {max_results_per_request} results. To access additional pages, issue a CONTINUE operation. The root query (id=0) cannot be continued. Think step-by-step and keep outputs structured for machine-readability.
[User]
<user_query>{user_query}</user_query>
<current_iteration>{current_iteration}</current_iteration>
<last_iteration_state>{last_iteration_state}</last_iteration_state>
<previous_iteration_state>{previous_iteration_state}</previous_iteration_state>
<last_checklist>{last_checklist}</last_checklist>
<last_experience_replay>{last_experience_replay}</last_experience_replay>
<all_subqueries>{all_subqueries}</all_subqueries>
<strategies>(See Table 12)</strategies>
<linking_guidance>(See Table 12)</linking_guidance>
<field_reference>(See Table 12)</field_reference>
<instructions>
1) Plan & Validate: Briefly reflect on research gaps and progress. Validate <last_iteration_state> against <last_checklist> (note any mismatches between requested target_k vs. actual retrieval counts).
2) Subqueries Strategy: CONTINUE, DERIVE, EXPAND.
3) Subqueries: Propose 3–6 keyword-style subqueries with integer target_k each. For each item, choose source_id from
<all_subqueries> (including 0 for DERIVE/EXPAND) and set link_type.
4) Checklist: Write concrete retrieval criteria tailored to the current subqueries (what to retrieve/include/exclude; methods, tasks, datasets, metrics, temporal filters).
5) Experience Replay: Update long-term memory by integrating
<last_experience_replay> with insights from <last_iteration_state> and current rationale. Preserve historical key points and explicitly note what changed vs. the previous iteration.
6) Completion: Set is_complete=true ONLY if existing retrieved information is sufficient to comprehensively answer the user query.
</instructions>
<output_format>
Provide JSON only inside <planner_output> tags:
<planner_output>
{"subqueries": [{"link_type": "...", "source_id": $x$, "text": "...", "target_k": xx}, ...],
"checklist": "...", "experience_replay": "...", "is_complete": false}
</planner_output>
</output_format>
Figure 8: Query Planning prompt template with full history and state validation.
Output Example:
  "subqueries": [
    {"link_type": "continue", "source_id": 3, "target_k": 10},
    {"link_type": "derive", "source_id": 4,
     "text": "self-normalized IPS SNIPS Wang 2017", "target_k": 10},
    {"link_type": "expand", "source_id": 0,
     "text": "IPS variance reduction bias correction", "target_k": 10}
  "checklist": "Prioritize IPS/SNIPS for selection bias. Target foundational
     SNIPS paper (Wang et al. 2017). Include theoretical analyses...",
  "experience_replay": "Iter 1: IPS/SNIPS concentrated in counterfactual
     learning. Foundational paper missing. Iter 2 derives targeted queries...",
  "is_complete": false
Figure 9: Output example from Query Planning.
E.2 Relevance Assessment Prompts
This stage evaluates candidates $\{\mathcal{C}_{i}\}$ against plan $\mathcal{P}_{t}$ and outputs selected papers $\mathcal{S}_{t}$ with feedback $\mathcal{O}_{t}$ (cf. Section 3.1). We implement two modes.
E.2.1 Abstract-Only Mode
Candidates are classified based on metadata alone. Figure 10 presents the template; Figure 11 shows a representative output.
Prompt: Relevance Assessment (Abstract-only) [System]
You are the Relevance Assessment module. Given a subquery, the Query Planning checklist, the assessment recipe, and candidate papers with retriever scores, decide which papers to maintain. Be strict; maintain only high-quality, directly relevant papers. Output decisions in JSON. Additionally, provide a structured overview of the retrieval and selection results to help the Query Planning module adjust future subqueries.
[User]
<original_query>{user_query}</original_query>
<sub_query>{sub_query}</sub_query>
<planner_checklist>{planner_checklist}</planner_checklist>
<selector_recipe>{selector_recipe}</selector_recipe>
<candidates>{candidates}</candidates>
<instructions>
Return JSON inside <selector_output>...</selector_output>. Fields:
- selected: array of paper_id kept
- reasons: mapping paper_id $\rightarrow$ short reason
- overview: a single string summarizing the result, including:
 1) retrieved_topics: what topics the retrieved papers cover
 2) relevant_summary: what the selected papers discuss
 3) irrelevant_summary: what discarded papers discuss and why irrelevant
 4) adjustment_suggestions: optional suggestions for refining the subquery
</instructions>
<output_format>
<selector_output>
{"selected": [], "reasons": {}, "overview": "..."}
</selector_output>
</output_format>
Figure 10: Relevance Assessment prompt template (Abstract-only mode).
Output Example:
  "selected": ["1503.02045"],
  "reasons": {
    "1503.02045": "Addresses estimation after parameter selection,
       discussing selection bias and corrective estimators."
  "overview": "1.retrieved_topics: Wireless comm, numerical methods,
     covariance selection. 2.relevant_summary: Selected paper tackles
     selection bias in parameter estimation. 3.irrelevant_summary: Most
     papers discuss ’selection’ in unrelated contexts. 4.adjustment:
     Use ’Inverse Propensity Score weighting selection bias’."
Figure 11: Output example from Relevance Assessment (Abstract-only).
E.2.2 Adaptive Browsing Mode
An intermediate to_browse label triggers full-text extraction before final classification. Figure 12 presents the template; Figure 13 shows a representative output.
Prompt: Relevance Assessment (Adaptive Browsing) [System]
You are the Relevance Assessment module. Given a subquery, the Query Planning checklist, the assessment recipe, and candidate papers (meta-info and/or browser evidence), decide which papers to select, discard, or browse further. If information is insufficient, delegate to the Browser. Be strict on selection; be precise on browsing goals. Output decisions in JSON. Additionally, provide a structured overview to help the Query Planning module adjust future subqueries.
[User]
<original_query>{user_query}</original_query>
<sub_query>{sub_query}</sub_query>
<planner_checklist>{planner_checklist}</planner_checklist>
<selector_recipe>{selector_recipe}</selector_recipe>
<context><candidates>{candidates}</candidates></context>
<instructions>
1.
Finalize Decisions (with Browser Evidence): For candidates that include browser_summary data, make a definitive decision: include them in selected or discarded. Do not put these candidates into to_browse.
2.
Handle Meta-info Only Candidates: For candidates with only meta-info (title, abstract…): if relevance is clear, add to selected or discarded; if relevance is plausible but uncertain, add to to_browse with a specific extraction goal.
3.
Generate Overview: Generate an overview based strictly on the current batch of candidates and decisions. Summarize insights gained from both selected papers and browser results to help the Query Planning module refine future subqueries.
4.
JSON Requirement: Output the result strictly in JSON format inside <selector_output>.
Fields (ensure all paper_id references use arxiv_id format):
- selected: array of paper_id kept as high-quality, relevant evidence
- discarded: array of paper_id excluded from further consideration
- to_browse: mapping paper_id $\rightarrow$ specific extraction goal for the Browser
- reasons: mapping paper_id $\rightarrow$ short reason for the current decision
- overview: a single string (retrieved_topics, relevant_summary, irrelevant_summary, adjustment_suggestions)
</instructions>
<output_format>
<selector_output>
{"selected": [], "discarded": [], "to_browse": {}, "reasons": {}, "overview": "..."}
</selector_output>
</output_format>
Figure 12: Relevance Assessment prompt template (Adaptive Browsing mode).
Output Example:
  "selected": ["1003.5956"],
  "discarded": ["1504.06937", "1310.1404", "1407.2806"],
  "to_browse": {},
  "reasons": {
    "1003.5956": "Seminal paper on unbiased offline evaluation,
       connects to IPS/SNIPS correction for selection bias.",
    "1504.06937": "Constrained bandits regret. No IPS/SNIPS."
  "overview": "1.retrieved_topics: Contextual bandits, cold-start.
     2.relevant_summary: One paper proposes unbiased offline evaluation.
     3.irrelevant_summary: Nine papers focus on online bandit algorithms.
     4.adjustment: Include ’unbiased offline evaluation’."
Figure 13: Output example from Relevance Assessment (Adaptive Browsing).
E.3 Browser Extraction Prompt
This module extracts targeted evidence from full-text when invoked by adaptive browsing. Figure 14 presents the template; Figure 15 shows a representative output.
Prompt: Browser Extraction [System]
You are the Paper Extraction module. Your goal is to extract the minimum viable evidence from a paper to satisfy a specific research goal. Focus on precision over volume. Capture technical verbatim data accurately but eliminate tangential context.
[User]
<paper_content>{full_text}</paper_content>
<user_goal>{task}</user_goal>
<instructions>
Return JSON inside <extractor_output>. All fields must be Strings:
- rational: List only the specific section names relevant to the goal.
- evidence: Extract concise, verbatim snippets that directly support the goal. Use ellipses (…) to skip irrelevant filler.
- summary: A max 3-sentence synthesis stating what the paper provides regarding the goal.
</instructions>
<output_format>
<extractor_output>
{"rational": "...", "evidence": "...", "summary": "..."}
</extractor_output>
</output_format>
Figure 14: Browser Extraction prompt template.
Output Example:
  "rational": "Section 2: Related Work, Section 7: References",
  "evidence": "Point-based rendering traces back to Levoy [1985]...
     Zwicker et al. [2001] introduced surface splatting...
     Yifan et al. [2019] proposed differentiable surface splatting...",
  "summary": "Survey traces Gaussian-based rendering lineage from
     Zwicker (2001) through Yifan (2019) to 3D Gaussian Splatting.
     Confirms historical connection between EWA and differentiable
     point-based graphics."
Figure 15: Output example from Browser Extraction.